OpenStudy (anonymous):

3x+4y=4 x=55-9y Solve by the substitution method.

OpenStudy (anonymous):

3(55-9y)+4y=4

OpenStudy (anonymous):

165-27y+4y=4

OpenStudy (anonymous):

165-23y=4

OpenStudy (anonymous):

-23y=-161

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

x=55-9(7) x=-8

OpenStudy (anonymous):

finished :d

OpenStudy (anonymous):

step by step